OpenStudy (anonymous):

Two identical objects each with a charge of +3.0 μC experience a repulsive force of +0.90 N between them. How far apart are the objects? A. 0.10 m B. 0.20 m C. 0.30 m D. 0.90 m

OpenStudy (michele_laino):

we have to apply the law of Coulomb \[F = K\frac{{{Q_1}{Q_2}}}{{{d^2}}}\]

OpenStudy (anonymous):

ok! what do we plug in? :/

OpenStudy (michele_laino):

so we can write: \[d = \sqrt {K\frac{{{Q_1}{Q_2}}}{F}} = \sqrt {9 \times {{10}^9}\frac{{3 \times {{10}^{ - 6}} \times 3 \times {{10}^{ - 6}}}}{{0.9}}} \]

OpenStudy (anonymous):

ok! and we get 0.3 ?

OpenStudy (anonymous):

and choice C is our solution?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay! thanks!!:)

OpenStudy (michele_laino):

:)